# Robust Blockwise Random Pivoting:

### v3

@@ 1.3 Notations and Roadmap
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, let $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}=\sum_{i=1}^{\min(n,d)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the singular value decomposition of $\mathbf{X}$ with $\sigma_{1}\geq\cdots\geq\sigma_{\min(n,d)}\geq 0$. Given any $r\in[\min(n,d)]$, we denote $\mathbf{U}_{r}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\bm{\Sigma}_{r}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{r}\right)$, $\mathbf{V}_{r}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$ such that ${\mathbf{X}}_{\langle r\rangle}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$.
We follow the MATLAB notation for vector and matrix indexing throughout this work. For any $n\in\mathbb{N}$, let $[n]=\left\{1,\cdots,n\right\}$ and $\Delta_{n}=\left\{\mathbf{p}\in[0,1]^{n}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{p}\right\|_{1}=1\right\}$ be the probability simplex of dimension $n$. Also, let $\mathfrak{S}_{n}$ be the set of all permutations of $[n]$. For any distribution $\mathbf{p}\in\Delta_{n}$ and $k\in\mathbb{N}$, $\mathbf{p}^{k}$ represents the joint distribution over $\Delta_{n}^{k}=\Delta_{n}\times\cdots\times\Delta_{n}$ such that $S=\left\{s_{j}\sim\mathbf{p}\right\}_{j\in[k]}\sim\mathbf{p}^{k}$ is a set of $k$ i.i.d. samples. For any $n\in\mathbb{N}$, let $\mathbf{e}_{i}$ ($i\in[n]$) be the $i$-th canonical base of $\mathbb{R}^{n}$. For any $m\in\mathbb{N}$, let $\textbf{1}_{m}\in\mathbb{R}^{m}$ and $\textbf{0}_{m}\in\mathbb{R}^{m}$ be the vector with all entries equal to one and zero, respectively.
As a roadmap, we review the adaptiveness-only and randomness-only skeleton selection algorithms in Section 2, along with the two different combinations of adaptiveness and randomness in Section 3. Then, we introduce the RBRP algorithm formally in Section 4. In Section 5, we discuss the efficient construction of an interpolation matrix for exactly and approximately ID-revealing algorithms. In Section 6, we demonstrate the appealing empirical accuracy and efficiency of RBRP via extensive numerical experiments. We provide the experiment code, as well as a high-performance implementation of RBRP, at https://github.com/dyjdongyijun/Robust_Blockwise_Random_Pivoting.
@@ 2.1 Adaptiveness via Greedy Squared-norm Pivoting
Greedy pivoting is a classical way of incorporating adaptiveness in skeleton selection [24, 44, 48]. Column pivoted QR (CPQR) [22, Section 5.4.2] is one of the most commonly used greedy pivoting methods. The pivoting strategy and adaptive update are two key components that characterize a greedy pivoting method. In particular, CPQR involves greedy squared-norm pivoting and adaptive QR updates as synopsized below.
Starting with $\mathbf{X}^{(0)}\leftarrow\mathbf{X}$, given a(n) (active) data (sub)matrix $\mathbf{X}^{(t)}\in\mathbb{R}^{n\times d}$ with at most $n-t$ nonzero rows, CPQR on $\mathbf{X}^{\top}$ greedily pivots the row in $\mathbf{X}^{(t)}$ with the maximum $\ell_{2}$-norm $s_{t+1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{X}^{(t)}\left(i,:\right)\right\|_{2}^{2}$ (i.e. squared-norm pivoting) and adaptively updates the active submatrix by projecting it onto the orthogonal complement of the selected skeleton222 For illustration, we consider the exact arithmetic here and express the QR update as a Gram-Schmidt process, whereas, in practice, QR updates are usually conducted via Householder reflection [22, Section 5.1.2] for numerical stability. :
(4) $\mathbf{X}^{(t+1)}\leftarrow\mathbf{X}^{(t)}-\mathbf{X}^{(t)}\frac{\mathbf{X}^{(t)}(s_{t+1},:)^{\top}\mathbf{X}^{(t)}(s_{t+1},:)}{\left\|\mathbf{X}^{(t)}(s_{t+1},:)\right\|_{2}^{2}}.$
The rank-revealing property of CPQR [24, Theorem 7.2] implies that the first $k$ pivots $S$ consisting of $\left\{s_{1},\cdots,s_{k}\right\}$ form a reasonable skeleton subset with
(5) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq 4^{k}\left(n-k\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle k\rangle}\right\|_{F}^{2}.$
The exponential dependence on $k$ in Equation 5 is tight under adversarial inputs (e.g. Kahan matrices [29][24, Example 1]). In the worst case, CPQR can take almost all of the $n$ points as skeletons before finding an $(r,\epsilon)$-ID. Concretely, [5, Theorem 4.1] implies that for CPQR,
(6) $\mathcal{E}_{\mathbf{X}}\left(S\right)\leq\frac{1}{\eta_{r}}\cdot\left(1-\frac{k}{n}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
That is, CPQR provides an $(r,\epsilon)$-ID when $k\geq\left(1-(1+\epsilon)\eta_{r}\right)n$. Moreover, for any $S$ with $\left|S\right|=k\leq\min\left(1-(1+\epsilon)\eta_{r},\eta_{r}\right)n-1$, there exists an $\mathbf{X}$ where CPQR fails to form an $(r,\epsilon)$-ID.
@@
Remark 2.2 (CPQR is inherently sequential but error-revealing).
From the computational efficiency perspective, a major drawback of CPQR is that the adaptive updates are inherently sequential (only half of the operations can be cast into high-performance matrix-matrix multiplication in the best-known algorithm that underlies LAPACK’s routine geqp3 [41]). As a result, CPQR tends to be observably slower than simple alternatives with the same asymptotic complexity like Gaussian elimination with partial pivoting [22, Section 3.4]—another common greedy pivoting method that retains parallelizability at a cost of the rank-revealing guarantee.
Nevertheless, the adaptive nature grants CPQR an appealing bonus—the error-revealing ability. It is known that the skeletonization error of CPQR, i.e. $\mathcal{E}_{\mathbf{X}}\left(S\right)=\left\|\mathbf{X}^{\left(\left|S\right|\right)}\right\|_{F}^{2}$, can be downdated efficiently at the end of each update.
@@ 2.2 Randomness via Squared-norm Sampling
Sampling is a widely used approach for random skeleton selection (also referred to as column subset selection, e.g. in [11, 9]). To identify a nearly optimal (smallest possible) skeleton subset $S$ for an $(r,\epsilon)$-ID, the intuitive goal is to construct a distribution $\mathbf{p}\in\Delta_{n}$ over the $n$ data points $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ that weights each $\mathbf{x}_{i}$ according to its relative “importance.”
Analogous to the squared-norm pivoting in CPQR (Remark 2.1), sampling according to the squared-norm distribution is a natural choice for such “importance” sampling. Concretely, adapting results from [13] and [5, Theorem 4.2] leads to the following expected skeleton complexity for squared-norm sampling.
Proposition 2.3 (Squared-norm sampling).
Consider a size-$k$ skeleton subset $S=\left\{s_{1},\cdots,s_{k}\right\}$ where each $s_{j}\in S$ is sampled from the squared-norm distribution $s_{j}\sim\mathbf{p}=\left(p_{1},\cdots,p_{n}\right)$, $p_{i}=\frac{\left\|\mathbf{x}_{i}\right\|_{2}^{2}}{\left\|\mathbf{X}\right\|_{F}^{2}}$ independently with replacement. Then
$\mathbb{E}_{S\sim\mathbf{p}^{k}}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\frac{r-1}{k\eta_{r}}+\frac{1}{k}\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
That is, squared-norm sampling yields an $(r,\epsilon)$-ID when $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$.
Compared to the skeleton complexity guarantee of squared-norm pivoting (CPQR) in Equation 6, the squared-norm sampling in Proposition 2.3 provides a potentially weaker guarantee in expectation (instead of worst-case) but yields a better skeleton complexity that scales proportionally to $r$ (instead of $n$).
Despite the improved expected skeleton complexity compared to CPQR, like CPQR (Remark 2.1), squared-norm sampling can also be vulnerable to adversarial inputs by selecting redundant skeletons with higher probabilities. Such potential inefficiency can be reflected by the linear dependence of the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ on $\frac{r}{\eta_{r}}$, as instantiated in Example 2.4.
@@
For $n\in\mathbb{N}$ divisible by $k\in\mathbb{N}$, $r\in\mathbb{N}$ such that $k=(1+\beta)r<n$ for some $\beta\in\mathbb{N}$, letting $\alpha_{i}=\sqrt{\alpha}\gg\beta\geq 1$ for all $1\leq i\leq r$, and $\alpha_{i}=1$ for all $r+1\leq i\leq k$, we consider the following data matrix
@@
Such inefficiency is mirrored in the skeleton complexity $k\geq\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ with a small $\eta_{r}$:
@@
It is worth highlighting that adaptiveness effectively remedies this inefficiency of squared-norm sampling. For instance, after picking $\alpha_{i}\mathbf{e}_{i}$ as the first skeleton, the adaptive update (e.g. Equation 4) eliminates the remaining $\frac{n}{k}-1$ rows of $\mathbf{X}$ in the $\mathbf{e}_{i}$ direction and therefore helps exclude the possible redundant samples.
As the well-studied alternatives to squared-norm sampling, determinantal point process (DPP sampling) [2, 12] and leverage score sampling [33] provide better skeleton complexity guarantees but with higher cost of constructing the corresponding distributions as a trade-off.
For example, the k-DPP sampling [31] is known for its nearly optimal skeleton complexity in expectation [2, 25, 5]:
(7) $k=\left|S\right|\geq\frac{r}{\epsilon}+r-1\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \mathbb{E}_{S\sim\text{k-DPP}(\mathbf{X})}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq(1+\epsilon)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}.$
As a trade-off, the classical SVD-based algorithm for k-DPP$(\mathbf{X})$ [27] takes $O\left(nd^{2}\right)$ to construct the distribution and $O\left(nk^{2}\right)$ to draw $k$ samples.
@@ 3 Combining Adaptiveness and Randomness
We recall that the vulnerability of greedy pivoting (Remark 2.1) comes from the scarce adversarial inputs which can be negligible in expectation under randomization. Meanwhile, the vulnerability of squared-norm sampling (Example 2.4) can be effectively circumvented by incorporating adaptive updates. Therefore, a natural question is how to combine adaptiveness and randomness effectively for better skeleton selection. In this section, we review two existing ideas that involve both adaptiveness and randomness in different ways, while comparing their respective advantages and drawbacks.
@@ 3.1 Random Pivoting
Random pivoting [13, 14, 1, 5] is arguably the most intuitive skeleton selection method that combines adaptiveness and randomness. It can be simply described as “replacing the greedy squared-norm pivoting in CPQR with squared-norm sampling” or “adaptively updating the dataset via QR after each step of squared-norm sampling,” as formalized in Algorithm 1.
Algorithm 1 Sequential random pivoting (SRP)
@@
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
@@
Algorithm 1 has an asymptotic complexity of $O\left(ndk+dk^{2}\right)$, with $k$ inherently sequential passes through $\mathbf{X}$ (or its submatrices), while the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively. Meanwhile, analogous to greedy pivoting (Remark 2.2), Algorithm 1 is error-revealing thanks to its adaptive nature.
@@
The idea of random pivoting in Algorithm 1 is ubiquitous in various related problems, e.g. the combination of volume sampling and adaptive sampling for low-rank approximation [14], the Randomly Pivoted Cholesky (RPCholesky) [5] for kernel column Nyström approximation [35, 19.2], and the $D^{2}$-sampling for k-means++ clustering [1].
Specifically, applying RPCholesky [5, Algorithm 2.1] to the kernel matrix $\mathbf{X}\mathbf{X}^{\top}\in\mathbb{R}^{n\times n}$ can be reformulated as Algorithm 2, which is equivalent to Algorithm 1 in the exact arithmetic (Remark 3.1). Compared to Algorithm 1, the skeleton selection stage of Algorithm 2 has an asymptotic complexity of $O\left(ndk+nk^{2}\right)$ (cf. $O\left(ndk+dk^{2}\right)$ for Algorithm 1), while only requiring $O(nk)$ memory for the storage of $\mathbf{L}^{(t)}$.
Algorithm 2 RPCholesky [5, Algorithm 2.1]
1:$\mathbf{L}^{(0)}\leftarrow\textbf{0}_{n\times 0}$, $\bm{\pi}\leftarrow[1,\cdots,n]$, $S\leftarrow\emptyset$, $t=0$
2:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
3:while $\left\|\mathbf{d}^{(t)}\right\|_{1}<\tau\left\|\mathbf{d}^{(0)}\right\|_{1}$ (or $\left|S\right|<k$do
4:     $t\leftarrow t+1$
5:     $s_{t}\sim\mathbf{d}^{(t-1)}/\left\|\mathbf{d}^{(t-1)}\right\|_{1}$
6:     $S\leftarrow S\cup\left\{s_{t}\right\}$, $\bm{\pi}\leftarrow\mathtt{swap}\left(\bm{\pi},t,s_{t}\right)$triangleright $\bm{\pi}(1:t)$ consists of $S$
7:     $\mathbf{a}^{(t)}\leftarrow\textbf{0}_{n}$, $\bm{\pi}_{a}\leftarrow\bm{\pi}(t:n)$triangleright “a” for active
8:     $\mathbf{a}^{(t)}\left(\bm{\pi}_{a}\right)\leftarrow\mathbf{X}\left(\bm{\pi}_{a},:\right)\mathbf{x}_{s_{t}}-\mathbf{L}^{(t-1)}\left(\bm{\pi}_{a},1:t-1\right){\mathbf{L}^{(t-1)}\left(s_{t},1:t-1\right)}^{\top}$triangleright $O\left(nd+nt\right)$
9:     $\mathbf{L}^{(t)}\leftarrow\left[\mathbf{L}^{(t-1)},\mathbf{a}^{(t)}/\sqrt{\mathbf{a}^{(t)}\left(s_{t}\right)}\right]$triangleright $\mathbf{a}^{(t)}\left(s_{t}\right)=\mathbf{d}^{(t-1)}(s_{t})$
10:     $\mathbf{d}^{(t)}(i)\leftarrow 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in S$, $\mathbf{d}^{(t)}(i)\leftarrow\mathbf{d}^{(t-1)}(i)-\frac{\left(\mathbf{a}^{(t)}(i)\right)^{2}}{\mathbf{a}^{(t)}\left(s_{t}\right)}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\notin S$triangleright $O(n)$
Remark 3.1 (Relation between QR with sequential random pivoting and RPCholesky).
Sequential random pivoting (Algorithm 1) on $\mathbf{X}$ is equivalent to RPCholesky (Algorithm 2 [5, Algorithm 2.1]) on $\mathbf{X}\mathbf{X}^{\top}$ in the exact arithmetic.
Proof 3.2 (Rationale for Remark 3.1).
To show the equivalence in the exact arithmetic, it is sufficient to observe that Algorithms 1 and 2 yield the same $\mathbf{a}^{(t)}$ and $\mathbf{d}^{(t)}$ in each iteration $t=1,2,\cdots$.
By induction, when $t=1$, both Algorithms 1 and 2 lead to $\mathbf{a}^{(1)}=\mathbf{X}\mathbf{x}_{s_{1}}$ and
$\mathbf{d}^{(1)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\mathbf{x}_{s_{1}}\right)^{2}}{\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}}=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(1)}\left(\mathbf{Q}^{(1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}\quad\forall\leavevmode\nobreak\ i\in[n]\setminus\left\{s_{1}\right\},$
while $\mathbf{d}^{(1)}(s_{1})=0$. We also observe that $\mathbf{L}^{(1)}=\mathbf{X}\mathbf{Q}^{(1)}$ since $\mathbf{a}^{(1)}\left(s_{1}\right)=\left\|\mathbf{x}_{s_{1}}\right\|_{2}^{2}$.
Suppose that in the $(t-1)$-th iteration, Algorithms 1 and 2 produce equivalent $\mathbf{a}^{(t-1)}$ and $\mathbf{d}^{(t-1)}$ under the exact arithmetic, while $\mathbf{d}^{(t-1)}\left(i\right)=\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}$ for all $i\in[n]$ and $\mathbf{L}^{(t-1)}=\mathbf{X}\mathbf{Q}^{(t-1)}$.
Then in the $t$-th iteration, both Algorithms 1 and 2 compute
$\mathbf{a}^{(t)}=\mathbf{X}\mathbf{x}_{s_{t}}-\mathbf{X}\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\mathbf{x}_{s_{t}}=\mathbf{X}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}},$
$\mathbf{L}^{(t)}=\mathbf{X}\mathbf{Q}^{(t)}$, and for all $i\in[n]$,
$\mathbf{d}^{(t)}(i)=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}-\frac{\left(\mathbf{x}_{i}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}\right)^{2}}{\mathbf{x}_{s_{t}}^{\top}\left(\mathbf{I}_{d}-\mathbf{Q}^{(t-1)}\left(\mathbf{Q}^{(t-1)}\right)^{\top}\right)\mathbf{x}_{s_{t}}}$
$=$ $\left\|\left(\mathbf{I}_{d}-\mathbf{Q}^{(t)}\left(\mathbf{Q}^{(t)}\right)^{\top}\right)\mathbf{x}_{i}\right\|_{2}^{2}.$
In light of the equivalence between Algorithms 1 and 2, [5, Theorem 3.1] provides the following skeleton complexity guarantee for sequential random pivoting.
Proposition 3.3 (Sequential random pivoting).

### v5

@@ 1.3 Notations and Roadmap
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, let $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}=\sum_{i=1}^{\min(n,d)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the singular value decomposition of $\mathbf{X}$ with $\sigma_{1}\geqslant\cdots\geqslant\sigma_{\min(n,d)}\geqslant 0$. Given any $r\in[\min(n,d)]$, we denote $\mathbf{U}_{r}=\left[\mathbf{u}_{1},\cdots,\mathbf{u}_{r}\right]$, $\bm{\Sigma}_{r}=\mathop{\mathrm{diag}}\left(\sigma_{1},\cdots,\sigma_{r}\right)$, $\mathbf{V}_{r}=\left[\mathbf{v}_{1},\cdots,\mathbf{v}_{r}\right]$ such that ${\langle\mathbf{X}\rangle}_{r}=\mathbf{U}_{r}\bm{\Sigma}_{r}\mathbf{V}_{r}^{\top}$. We follow the MATLAB notation for vector and matrix indexing throughout this work. For any $n\in\mathbb{N}$, let $[n]=\left\{1,\cdots,n\right\}$ and $\Delta_{n}\in[0,1]^{n}$ be the probability simplex of dimension $n$. Also, let $\mathfrak{S}_{n}$ be the set of all permutations of $[n]$. For any distribution $\mathbf{p}\in\Delta_{n}$ and $k\in\mathbb{N}$, $\mathbf{p}^{k}$ represents the joint distribution over $\Delta_{n}^{k}=\Delta_{n}\times\cdots\times\Delta_{n}$ such that $S=\left\{s_{j}\sim\mathbf{p}\right\}_{j\in[k]}\sim\mathbf{p}^{k}$ is a set of $k$ i.i.d. samples. For any $n\in\mathbb{N}$, let $\mathbf{e}_{i}$ ($i\in[n]$) be the $i$-th canonical base of $\mathbb{R}^{n}$. For any $m\in\mathbb{N}$, let $\textbf{1}_{m}\in\mathbb{R}^{m}$ and $\textbf{0}_{m}\in\mathbb{R}^{m}$ be the vector with all entries equal to one and zero, respectively. We adapt the standard asymptotic notations: for any functions $f,g:\mathbb{R}_{+}\to\mathbb{R}_{+}$, we write $f=O\left(g\right)$ or $f\lesssim g$ if there exists some constant $C>0$ such that $f(x)\leqslant Cg(x)$ for all $x\in\mathbb{R}_{+}$; $f=\Omega\left(g\right)$ or $f\gtrsim g$ if $g=O\left(f\right)$; $f=\Theta(g)$ or $f\asymp g$ if $f=O\left(g\right)$ and $f=\Omega\left(g\right)$.
As a roadmap, we review the adaptiveness-only and randomness-only skeleton selection algorithms in Section 2, along with the two different combinations of adaptiveness and randomness in Section 3. Then, we introduce the RBRP algorithm formally in Section 4. In Section 5, we discuss the efficient construction of interpolation matrices after skeleton selection. In Section 6, we demonstrate the appealing empirical accuracy and efficiency of RBRP via extensive numerical experiments. The code, as well as a high-performance implementation of RBRP, is available at https://github.com/dyjdongyijun/Robust_Blockwise_Random_Pivoting.
@@ 2.1 Adaptiveness via Greedy Squared-norm Pivoting
Greedy pivoting is a classical way of incorporating adaptiveness in skeleton selection [35, 60, 62]. Column pivoted QR (CPQR) [33, Section 5.4.2] is one of the most commonly used greedy pivoting methods. The pivoting strategy and adaptive update are two key components that characterize a greedy pivoting method. In particular, given an active submatrix $\mathbf{X}^{(t)}\in\mathbb{R}^{n\times d}$ with at most $n-t$ nonzero rows, CPQR involves greedy squared-norm pivoting and adaptive QR updates222 Notice that (4) can be viewed as the $(t+1)$th-step of either CPQR on $\mathbf{X}^{\top}$ or row pivoted LQ (RPLQ) on $\mathbf{X}$. Throughout this work, we refer to column pivoting on $\mathbf{X}^{\top}$ as CPQR without ambiguity. In (4), we consider the exact arithmetic and express the QR update as a Gram-Schmidt process for illustration purposes. In practice, QR updates are usually conducted via Householder reflection [33, Section 5.1.2] for numerical stability. :
(4) $s_{t+1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{X}^{(t)}\left(i,:\right)\right\|_{2}^{2},\ \mathbf{X}^{(t+1)}\leftarrow\mathbf{X}^{(t)}-\mathbf{X}^{(t)}\frac{\mathbf{X}^{(t)}(s_{t+1},:)^{\top}\mathbf{X}^{(t)}(s_{t+1},:)}{\left\|\mathbf{X}^{(t)}(s_{t+1},:)\right\|_{2}^{2}}.$
The rank-revealing property of CPQR [35, Theorem 7.2] implies that the first $k$ pivots $S=\left\{s_{1},\cdots,s_{k}\right\}$ form a reasonable skeleton subset with $\mathcal{E}_{\mathbf{X}}\left(S\right)\leqslant 4^{k}\left(n-k\right)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{k}\right\|_{F}^{2}$, where the exponential dependence on $k$ is tight under adversarial inputs (e.g. Kahan matrices [43][35, Example 1]). In the worst case, [8, Theorem C.2] shows that CPQR requires almost all the $n$ points ($k\geqslant\left(1-(1+\epsilon)\eta_{r}\right)n$) before finding a $(r,\epsilon)$-ID.
@@
Remark 2.2 (CPQR is inherently sequential but rank-adaptive).
From the efficiency perspective, a major drawback of CPQR is that the adaptive updates are inherently sequential (only half of the operations can be cast into high-performance matrix-matrix multiplication in the best-known algorithm that underlies LAPACK’s routine geqp3 [58]). As a result, CPQR tends to be observably slower than simple alternatives with the same asymptotic complexity like Gaussian elimination with partial pivoting [33, Section 3.4]—another common greedy pivoting method that retains parallelizability at a cost of the rank-revealing guarantee.
Nevertheless, the adaptive nature grants CPQR an appealing bonus—the rank-adaptiveness. It is known that the skeletonization error of CPQR, i.e. $\mathcal{E}_{\mathbf{X}}\left(S\right)=\left\|\mathbf{X}^{\left(\left|S\right|\right)}\right\|_{F}^{2}$, can be downdated efficiently at the end of each update.
@@ 2.2 Randomness via Squared-norm Sampling
Sampling is a widely used approach for random skeleton selection (also referred to as column subset selection, e.g. in [17, 14]), where the intuitive goal is to construct a distribution $\mathbf{p}\in\Delta_{n}$ over the $n$ data points $\left\{\mathbf{x}_{i}\right\}_{i\in[n]}$ according to their relative “importance.” Analogous to the squared-norm pivoting in CPQR ((4)), sampling according to the squared-norm distribution:
(5) $p_{i}={\left\|\mathbf{x}_{i}\right\|_{2}^{2}}/{\left\|\mathbf{X}\right\|_{F}^{2}}\quad\forall\leavevmode\nobreak\ i\in[n]$
is a natural choice for such “importance” sampling. In particular, [20] and [8, Theorem C.3] shows that squared-norm sampling requires $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ skeleton points to form a $(r,\epsilon)$-ID in expectation. Compared to the skeleton complexity of pivoting (CPQR), the squared-norm sampling provides a lower skeleton complexity in expectation that scales proportionally to $r$ instead of $n$.
Despite the improved expected skeleton complexity compared to CPQR, like CPQR, squared-norm sampling can also be vulnerable to adversarial inputs by selecting redundant skeletons with higher probabilities. Such potential inefficiency can be reflected by the linear dependence of the skeleton complexity $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ on $\frac{r}{\eta_{r}}$, as instantiated in Example 2.3.
@@
For $n\in\mathbb{N}$ divisible by $k\in\mathbb{N}$, $r\in\mathbb{N}$ such that $k=(1+\beta)r<n$ for some $\beta\in\mathbb{N}$, letting $\alpha_{i}=\sqrt{\alpha}\gg\beta\geqslant 1$ for all $1\leqslant i\leqslant r$, and $\alpha_{i}=1$ for all $r+1\leqslant i\leqslant k$, we consider the following data matrix
@@
Such inefficiency is mirrored in the skeleton complexity $k\geqslant\frac{r-1}{\epsilon\eta_{r}}+\frac{1}{\epsilon}$ with a small $\eta_{r}$:
@@
It is worth highlighting that adaptiveness effectively remedies this inefficiency of squared-norm sampling. For instance, after picking $\alpha_{i}\mathbf{e}_{i}$ as the first skeleton, the adaptive update (e.g. (4)) eliminates the remaining $\frac{n}{k}-1$ rows of $\mathbf{X}$ in the $\mathbf{e}_{i}$ direction and therefore helps exclude the possible redundant samples.
Determinantal point process (DPP) [5, 19] and leverage score sampling [49] are well-studied alternatives to squared-norm sampling that provide better skeleton complexities but with much higher cost of constructing the corresponding distributions as a trade-off. For example, the k-DPP sampling [45] is known for its nearly optimal skeleton complexity: as few as $k\geqslant\frac{r}{\epsilon}+r-1$ skeletons are sufficient to provide a $(r,\epsilon)$-ID in expectation [5, 36, 8]. As a trade-off, the classical SVD-based algorithm for k-DPP$(\mathbf{X})$ [6] takes $O\left(nd^{2}\right)$ to construct the distribution and $O\left(nk^{2}\right)$ to draw $k$ samples.
@@ 3 Combining Adaptiveness and Randomness
Recall that the vulnerability of greedy pivoting (Remark 2.1) comes from the scarce adversarial inputs. Under randomization, the probability that greedy pivoting keeps making adversarial skeleton selections becomes negligible. Meanwhile, the vulnerability of squared-norm sampling (Example 2.3) can be effectively circumvented by incorporating adaptive updates. Therefore, a natural question is how to combine adaptiveness and randomness effectively for better skeleton selection. In this section, we review two existing ideas that involve both adaptiveness and randomness in different ways, while comparing their respective advantages and drawbacks.
@@ 3.1 Random Pivoting
Random pivoting [20, 21, 2, 8] is arguably the most intuitive skeleton selection method that combines adaptiveness and randomness. It can be simply described as “replacing the greedy squared-norm pivoting in CPQR with squared-norm sampling” as formalized in Algorithm 1, with an asymptotic complexity of $O(ndk)$, consisting of $k$ inherently sequential passes through $\mathbf{X}$ (or its submatrices). The storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory. Analogous to greedy pivoting (Remark 2.2), Algorithm 1 is rank-adaptive thanks to its adaptive nature.
Algorithm 1 Sequential random pivoting (SRP) [8, Algorithm 7]
@@
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geqslant 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
@@
The idea of random pivoting is ubiquitous in various related problems, e.g. the combination of volume sampling and adaptive sampling for low-rank approximation [21], the Randomly Pivoted Cholesky (RPCholesky) [8] for kernel column Nyström approximation [52, 19.2], and the $D^{2}$-sampling for k-means++ clustering [2]. Specifically, in light of the close connection between partially pivoted QR and pivoted Cholesky factorization (see e.g. [38, §5.2] and [54, 8]), applying RPCholesky to $\mathbf{X}\mathbf{X}^{\top}\in\mathbb{R}^{n\times n}$ is equivalent to Algorithm 1 in exact arithmetic. In light of such equivalence, [8, Theorem 5.1] provides a nearly optimal skeleton complexity guarantee for sequential random pivoting:
Proposition 3.1 (Sequential random pivoting [8]).